determine the slope of the line that goes through each pair
(0,4) and (5,0)

Answer by stanbon(75887)

determine the slope of the line that goes through each pair
(0,4) and (5,0)
----------------
slope = change in y/ change in x = (0-4)/(5-0) = -4/5
=============================
Cheers,
stan H.